OpenStudy (anonymous):

What is the integral of tan(9x))??

OpenStudy (anonymous):

What is the area between the curves?? I really just need help with the integration part mostly.

OpenStudy (anonymous):

\[\int\limits_{-0.1}^{0.1} (2\sin9x) - (\tan(9x))\]

OpenStudy (anonymous):

\[\frac{ -2os(9x)) }{ 9 } - \int\limits_{?}^{?}\tan(9(x)) \]

OpenStudy (anonymous):

Thats what i have so far. IDK the integral of the tan9(x)). Are my bounds correct?

OpenStudy (anonymous):

integral of tan is funny. integral of tan(x) is -ln(cosx). knowing that, the integral of tan(9x) is -ln(cos9x)/9

OpenStudy (anonymous):

are we supposed to just memorize that b/c where did the ln come from?

OpenStudy (anonymous):

and the cos(x) part also??

OpenStudy (anonymous):

you can derive the integral of tan(x) by rewriting it as sin(x)cos(x)^-1, then use u substitution or whatever method you want. but briefly, since you add 1 to the exponent, the cos(x)^-1 becomes cos(x)^0, which is impossible and must be rewritten to ln(cos(x)).

OpenStudy (anonymous):

Oh ok i see now. So now have u looked at the file attached? Are my bounds correct?

OpenStudy (anonymous):

how will i subsititute -0.1 and 0.1 for x? Since cos(.9) cant e simplified

OpenStudy (anonymous):

I needed to find the integral of tan(9x) to find the area btwn the curves

OpenStudy (anonymous):

yes i believe your bounds are correct. um the integral of tan(9x) is -ln(cos9x)/9, plugging decimals into a log requires a calculator so that might be a problem for me..

OpenStudy (anonymous):

ok. thanks. I can do that part with a calculator. I just wanted to make sure my integrand and bounds were correct